OpenStudy (anonymous):

A certain radioactive substance has a half-life of 14 days. How many days will it take until there is only 12.5% of the original amount remaining?

OpenStudy (anonymous):

@amistre64 can u help me?

OpenStudy (amistre64):

i can

OpenStudy (amistre64):

we need to determine the rate .... and that can be done generically with some algebra

OpenStudy (amistre64):

spose we start with P and we end up with P/2 in a given amount of time \[P/2=Pe^{rt}\]divide of the P \[1/2=e^{rt}\]ln to undo the e \[ln(1/2)=rt\]divide off the t to solve for r \[\frac{ln(1/2)}{t}=r\]and since we know t is 14 days \[\frac{ln(1/2)}{14}=r\]

OpenStudy (amistre64):

now, spose we start with 100 parts and want to know when we will end up with 12.5 parts; we need to solve for t this time \[12.5 = 100e^{rt}\] work it thru and get to the end , plug in our r value we found to keep the clutter down

OpenStudy (amistre64):

.125 = e^rt ln(.125) = rt ln(.125) ------- = t r

OpenStudy (anonymous):

so its 42